4.5 grams of metallic chloride contains 35.5 gram of chlorine.the equivalent mass of metal is
a) 19.5         b)35.5 
c)39             d)78

Grade:12

2 Answers

Arun
4 years ago
Dear Amit
 
It should be 74.5 gm instead of 4.5 gm
mass of metal chloride= 74.5 g
mass of chlorine = 35.5 g
Thus mass of metal = mass of metal chloride - mass of chlorine= (74.5-35.5)= 39 g  
Equivalent weight is the weight of substance which reacts with 35.5 gram of chlorine.
Thus 39 gram of metal reacts with 35.5 g of chlorine.
Thus equivalent weight of metal is 39 gram.
Khimraj
4 years ago
Given : mass of metal chloride= 74.5 g
mass of chlorine = 35.5 g
Thus mass of metal = mass of metal chloride - mass of chlorine= (74.5-35.5)= 39 g  
Equivalent weight is the weight of substance which reacts with 35.5 gram of chlorine.
Thus 39 gram of metal reacts with 35.5 g of chlorine.
Thus equivalent weight of metal is 39 gram.